# Ivona Najdenkoska    Mohammad Mahdi Derakhshani11footnotemark: 1    Yuki M. Asano Nanne van Noord    Marcel Worring    Cees G. M. Snoek22footnotemark: 2

### v1

@@ 4.3 Ablation Study
Different types of Relative Positional Encodings In this experiment, we compare different types of relative positional encodings. We choose a recently introduced Contextual Position Encoding (CoPE) (Golovneva et al., 2024) for implementing the relative positional encodings in the text encoder. Afterwards, we perform the distillation and context length extension phases using the same dataset and parameters as with RoPE (Su et al., 2024). In Table 3 we observe that RoPE outperforms CoPE in long-caption retrieval tasks. This is due to RoPE’s strong ability to generalize across varying or extended sentence lengths, even beyond those on which the model was originally trained. In contrast, CoPE struggles to generalize as effectively when sequence lengths increase, as its embeddings are more dependent on the specific context within which they were trained. This explains why CoPE performs similarly to RoPE on the ShareGPT4V test split (same training distribution), but shows a larger performance gap on the Long-DCI and Urban-1k datasets.
@@
Table 3: Ablation comparing different relative position encodings: RoPE and CoPE in TULIP. RoPE generalizes better across varying or extended sentence lengths, especially on out-of-distribution datasets, namely Long-DCI and Urban-1k.
@@
Benefit of using cosine distillation loss. In this section, we vary the distillation loss function used during the relative position adaptation phase and report the results in Table 4. As seen, the cosine loss performs better than other alternatives across different datasets and tasks. We believe this is attributed to its alignment with the normalized embedding space of CLIP models and its scale invariance property. The scale invariance of cosine loss proves particularly beneficial in this distillation setup, where the student model (ViT-L-14 with the added RoPE or CoPE layer) can produce embeddings with different magnitudes compared to the teacher model. This invariance allows the distillation process to focus on transferring the essential directional information of the embeddings, rather than being influenced by potential scale discrepancies introduced by the RoPE or CoPE layer. As a result, cosine loss consistently outperforms other loss functions like L2 and MSE across various datasets (Long-DCI, ShareGPT4V, Urban-1K) and cross-modal tasks such as image-to-text and text-to-image, demonstrating its effectiveness in preserving the semantic relationships learned by the teacher model.
@@
Table 4: Ablation comparing different distillation loss terms in TULIP. Cosine loss yields the best performance across different datasets and tasks.
@@ Appendix A Appendix
A.1 Additional text-to-image generation results
Figure 7: Text-to-Image Generation results.
Figure 8: Text-to-Image Generation results.
Figure 9: Text-to-Image Generation results.

### v2

@@ 4.3 Ablation Study
Different types of positional encodings. In this experiment, we compare different types of positional encodings, namely absolute and relative ones. We choose a recently introduced Contextual Position Encoding (CoPE) (Golovneva et al., 2024) for implementing the relative positional encodings in the text encoder. Afterwards, we perform the distillation and context length extension phases using the same dataset and parameters as with RoPE (Su et al., 2024). In Table 3 we observe that RoPE outperforms CoPE in long-caption retrieval tasks. This is due to RoPE’s strong ability to generalize across varying or extended sentence lengths, even beyond those on which the model was originally trained. In contrast, CoPE struggles to generalize as effectively when sequence lengths increase, as its embeddings are more dependent on the specific context within which they were trained. This explains why CoPE performs similarly to RoPE on the ShareGPT4V test split (same training distribution), but shows a larger performance gap on the Long-DCI and Urban-1k datasets.
@@
Table 3: Ablation comparing different positional encodings such as Absolute, RoPE, and CoPE in TULIP. RoPE generalizes better across varying or extended sentence lengths, especially on out-of-distribution datasets, namely Long-DCI and Urban-1k.
@@
Benefit of using cosine distillation loss. In this section, we vary the distillation loss function used during the relative position adaptation phase and report the results in Table 4. As seen, the cosine loss performs better than other alternatives across different datasets and tasks. We believe this is attributed to its alignment with the normalized embedding space of CLIP models and its scale invariance property. The scale invariance of cosine loss proves particularly beneficial in this distillation setup, where the student model (ViT-L/14 with the added RoPE or CoPE layer) can produce embeddings with different magnitudes compared to the teacher model. This invariance allows the distillation process to focus on transferring the essential directional information of the embeddings, rather than being influenced by potential scale discrepancies introduced by the RoPE or CoPE layer. As a result, cosine loss consistently outperforms other loss functions like L2 and MSE across various datasets (Long-DCI, ShareGPT4V, Urban-1K) and cross-modal tasks such as image-to-text and text-to-image, demonstrating its effectiveness in preserving the semantic relationships learned by the teacher model.
@@
Table 4: Ablation comparing different distillation loss terms in TULIP. Cosine loss yields the best performance across different datasets and tasks. Note that here we report the performance of the distilled models before the relative position expansion phase.
@@ Appendix A Appendix
A.1 Qualitative evaluation of image generation using T5-based models, CLIP, Long-CLIP, and TULIP
Figure 7: Qualitative results of image generation using text encoders based on T5 (PIXART-Alpha and ELLA), CLIP, Long-CLIP, and TULIP. Our TULIP-based model demonstrates the ability to generate images with nuanced details from captions, offering a plug-and-play solution without requiring costly retraining.
Figure 8: Qualitative results of image generation using text encoders based on T5 (PIXART-Alpha and ELLA), CLIP, Long-CLIP, and TULIP. Our TULIP-based model demonstrates the ability to generate images with nuanced details from captions, offering a plug-and-play solution without requiring costly retraining.
A.2 Human evaluation of text-to-image generation
We design a human evaluation study by manually crafting captions and generating images by using Stable Diffusion XL (SDXL) with CLIP and TULIP-based text encoders. The resulting images are randomized, and annotators are tasked with selecting the image that is best aligned with the prompt (See Figure 9). They were instructed to evaluate alignment based on objects, their attributes, and their relationships. A total of 15 annotators reviewed 20 samples. We report the results in Table 5. As can be seen, the win ratios are 89% for the SDXL + TULIP model and 11% for the SDXL + CLIP model, reflecting the enhanced alignment capability that we observe in the qualitative comparisons as well. The following instructions were provided to the annotators:
Instructions for annotators: You are provided with 20 samples, each containing a prompt and two images generated by two different Stable Diffusion models. Your task is to compare the two images given the input prompt and choose the one you find more aligned with the prompt. Note that the differences might be very nuanced in some cases. Pay attention to the object attributes, relationships, verbs, object counting, etc.
Method Win rate (%)
SDXL + CLIP 11
SDXL + TULIP (Ours) 89
Table 5: Human evaluation results. The results reflect the enhanced alignment capability that we observe in the qualitative comparisons as well.
Figure 9: Examples used in the human evaluation study. We provide the text prompt and two randomized images generated by SDXL with different variants of the text encoder (TULIP vs CLIP).
A.3 Evaluation of compositional understanding benchmarks
CLIP-like models often struggle with accurately interpreting the compositional aspects of images, including spatial relationships, numeracy, and complex inter-object attributes. To evaluate TULIP’s capabilities in these areas, we considered two well-established baselines for multimodal compositional understanding: ARO (Yuksekgonul et al., 2023) and VL-Checklist (Zhao et al., 2022). The results in Table 6 indicate that TULIP consistently outperforms CLIP and surpasses Long-CLIP in 4 out of 7 settings. These findings highlight TULIP’s enhanced capability in handling fine-grained captions.
ARO VL-Checklist
Method VGR VGA Flickr COCO Obj Att Rel
CLIP 59.9 63.1 60.2 47.9 81.1 67.6 61.9
Long-CLIP 64.6 66.6 24.8 23.3 84.3 71.6 63.5
TULIP (Ours) 63.4 66.2 52.3 43.9 85.2 74.3 62.7
Table 6: Comparison across compositional understanding benchmarks, namely ARO and VL-Checklist. We can observe that TULIP consistently outperforms CLIP and surpasses Long-CLIP in 4 out of 7 settings, demonstrating its ability to handle fine-grained captions. Note that we use CLIP-ViT-L-14 as vision encoder.
A.4 DOCCI benchmark evaluation of long-caption retrieval
To additionally evaluate TULIP on long-caption retrieval tasks, we use DOCCI Onoe et al. (2024), which is a novel dataset consisting of images paired with long, human-annotated English descriptions. We performed experiments to compare our TULIP model to the baselines CLIP and Long-CLIP. We provide the top R@1 results in Table 7, where we observe that TULIP obtains the best performance, especially for image-to-text retrieval, demonstrating its efficiency and applicability across one more long-caption benchmark.
Img2Txt R@1 Txt2Img R@1
CLIP 65.7 63.1
Long-CLIP 66.5 78.5
TULIP (Ours) 77.9 79.1
Table 7: Long caption cross-modal retrieval comparison on DOCCI dataset. We observe that TULIP obtains the best performance, especially for image-to-text retrieval, demonstrating its efficiency and applicability across one more long-caption benchmark.
A.5 Training data amount for TULIP student model
To analyze the amount of data needed to train the student model, we conduct experiments using 33%, 66%, and the entire ShareGPT4V dataset (1.2 million image-caption pairs). The results on the retrieval tasks (top R@1) shown in Table 8, emphasize the importance of utilizing the full dataset, particularly for improving the performance on out-of-distribution data i.e. Long-DCI and Urban1k.
Long-DCI ShareGPT4V Urban-1K
Amount of data Img2Txt Txt2Img Img2Txt Txt2Img Img2Txt Txt2Img
33% 15.2 12.2 88.6 90.0 33.6 36.9
66% 28.7 26.2 94.6 95.5 37.6 45.0
100% 55.7 56.4 99.0 99.0 90.1 91.1
Table 8: Performance comparison across different dataset sizes.
A.6 Additional image generation comparison between TULIP and T5-based models
The T5 text encoder (Roberts et al., 2019) is well-known for its ability to handle long captions. To evaluate how our TULIP method compares to T5-based approaches, we conducted additional text-to-image generation experiments. In particular, we provide a qualitative comparison between our TULIP-based model and two T5-based models, PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024), in Figures 10, 11, 12, 13, 14 and 15. This comparison demonstrates that all models perform on par in generating high-quality images from long and nuanced text descriptions. However, our TULIP approach stands out due to its plug-and-play capability for image generation, requiring no additional fine-tuning, unlike PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024). This characteristic makes TULIP more versatile and practical compared to the fine-tuning demands of the T5-based counterparts.
Figure 10: Comparison to T5-based models (PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024)) for image generation.
Figure 11: Comparison to T5-based models (PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024)) for image generation.
Figure 12: Comparison to T5-based models (PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024)) for image generation.
Figure 13: Comparison to T5-based models (PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024)) for image generation.
Figure 14: Comparison to T5-based models (PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024)) for image generation.
Figure 15: Comparison to T5-based models (PIXART-Alpha (Chen et al., 2023a) and ELLA (Hu et al., 2024)) for image generation.
A.7 Qualitative evaluation of cross-modal retrieval using CLIP and TULIP
Figures 16 and 17 showcase selected results from our qualitative evaluation of cross-modal retrieval tasks. We can observe that our TULIP model excels at capturing fine-grained details in both captions and images, outperforming the baseline CLIP.
Figure 16: Qualitative comparison of image-to-text retrieval between CLIP and TULIP. We can observe that our TULIP model can capture the fine-grained details in both captions and images. Note that the red text color indicates visually misclassified concepts in the images, whereas the green color indicates correct ones.
Figure 17: Qualitative comparison of text-to-image retrieval between CLIP and TULIP. We can observe that our TULIP model can capture the fine-grained details in both captions and images. Note that the red text color indicates visually misclassified concepts in the images, whereas the green color indicates correct ones.